OpenStudy (arianna1453):

For what values of x is the graph of y= 2/5-x concave upward?

OpenStudy (arianna1453):



OpenStudy (arianna1453):

@TrojanPoem

OpenStudy (arianna1453):

@kaylaprincess

OpenStudy (trojanpoem):

y = 2(5-x)^-1 y' = -2(x-5)^-2 * -1 = 2(x-5)^-2 y'' = -4(x-5)^-3 when y'' is undefined x-5 = 0 x = 5